OpenStudy (anonymous):

I am putting oil,water and milk in microwave for 1 min. having same initial temperature and volume.Which will have Maximum Final Temperature and why? After what time temperature of oil exceed milk and water?

OpenStudy (anonymous):

first you'll have to know about the power output of the microwave. Is this a structured question?

OpenStudy (anonymous):

no...i have project to do.. power output of microwave=600 W

OpenStudy (anonymous):

from the eq \(Q=cm(t_1-t_0)\), and \(P=Q/tT\), \(PT=cm(t_1-t_0)\) where c is the specific heat capacity. Did you look up on the specific heat capacity on the objects? A higher heat capacity means lower final temperature.

OpenStudy (anonymous):

@shadowys a typo in the second equation -- the denominator is just 'T" and not 'tT'.

OpenStudy (anonymous):

right. I guess I forgot to delete the t